# deCIFer: Crystal Structure Prediction from Powder Diffraction Data using Autoregressive Language Models

### v2

@@ Abstract
deCIFer has two learnable components, $f_{{\mathbf{\Phi}}}$ and $f_{{\mathbf{\Theta}}}$. The encoder $f_{{\mathbf{\Phi}}}$ is a 2-layer MLP that takes a PXRD profile of dimension $1000$ and outputs a $512$-dimensional embedding. This embedding is prepended as the conditioning-token to the sequence of CIF tokens, where each CIF token is also of $D=512$ dimensions. $f_{{\mathbf{\Theta}}}$ is a decoder-only transformer Vaswani et al. (2017) with $8$ layers, each containing $8$ attention heads. We use a context length of $3076$ and a batch size of $32$. A linear warm-up of the learning rate is applied over the first $100$ epochs, followed by a cosine decay schedule using AdamW Loshchilov and Hutter (2017) ($\beta_{1}=0.9,\beta_{2}=0.95$, weight decay $10^{-1}$) for $50$K epochs with gradient accumulation of $40$ steps on a single NVIDIA A100 GPU with mixed-precision acceleration. All aspects of the model architecture were implemented in Pytorch Paszke et al. (2019). $f_{{\mathbf{\Phi}}}$ has $\approx\!0.78$M trainable parameters and $f_{{\mathbf{\Theta}}}$ contains $\approx\!26.94$M, resulting in a deCIFer model with $27.72$M parameters. The code is open-source and additional details can be found in Section A.8 in the Appendix.
@@
We conducted a series of experiments to evaluate deCIFer’s ability to perform CSP when conditioned on PXRD profiles, space group, and composition. First, we compare the baseline performance of deCIFer against an unconditioned variant (U-deCIFer) to assess the direct impact of PXRD conditioning on the generated structures. Next, we introduce controlled noise and peak broadening into the input PXRD data, examining deCIFer’s robustness under more challenging scenarios resembling real-world PXRD data. Finally, we apply deCIFer (trained on NOMA) to CHILI-100K Friis-Jensen et al. (2024) to demonstrate its scalability and performance on more complex crystal systems. In these experiments, we generate one CIF for each reference sample in the test sets. For consistent evaluation, each reference CIF is also processed through the fixed, clean transformation $\tau_{0}$ (FWHM $=0.05$, $\sigma_{\mathrm{noise}}^{2}=0$).
@@
Section 4.1 Importance of PXRD Conditioning
@@
Table 1 shows that deCIFer achieves lower average $R_{\mathrm{wp}}$ and WD than U-deCIFer across all descriptor settings. Figure 4 visualizes the distributions of $R_{\mathrm{wp}}$ and WD for both models, highlighting the large performance improvement due to PXRD conditioning, and a smaller but still notable improvement due to composition conditioning. Figure 5 shows that structures of more common crystal systems lead to better outcomes, while under-represented, lower symmetry crystal systems remain challenging. Figure 6 further illustrates the effectiveness of PXRD conditioning with three examples from the NOMA test set, showcasing the range of generated structures, from near-perfect alignment to structural mismatch. While U-deCIFer benefits from the addition of crystal descriptors, it never reaches the accuracy of deCIFer. Both models achieve high rates of syntactic validity, with the unconditional model marginally surpassing deCIFer (Table 2). Overall, these results demonstrate that conditioning on PXRD data and crystal descriptors substantially enhances the model’s ability to generate CIFs with close alignment to desired PXRD profiles.
Figure 4: Distribution of $R_{\mathrm{wp}}$ and WD for deCIFer and U-deCIFer on the NOMA test set, presented as violin plots with overlain boxplots; the median is shown for each distribution. Lower $R_{\mathrm{wp}}$ and WD values indicate a better alignment between the reference- and generated CIFs in terms of the reference PXRD profile. The models are evaluated in three settings: “none” (no crystal descriptors), “comp.” (composition only), and “comp. + s.g.” (composition plus space group).
Figure 5: Average metric values by crystal systems for deCIFer on the NOMA test set show better performance for well-represented systems, while rarer, low-symmetry systems lead to worse performance ($\uparrow R_{\mathrm{wp}},\uparrow$WD), reflecting their modelling complexity. The right-most plot shows crystal system distribution of the test set.
Figure 6: Three examples from the NOMA test set showcasing deCIFer generations. PXRD data is used for conditioning, with composition as a crystal descriptor. For each example, 100 CIFs are generated, and the median structure (based on $R_{\mathrm{wp}}$) is shown alongside the reference structure and PXRD profiles (noise-free, FWHM = $0.05$). The examples illustrate: (top) near-perfect PXRD conformity, (middle) structural match with suboptimal PXRD alignment, and (bottom) structural mismatch.
Table 1: Performance comparison for $20$K samples from the NOMA test set using deCIFer and U-deCIFer with descriptors. none: No crystal descriptors, comp.: compositional information only, and comp.+ s.g.: both compositional and space group information.
Desc. Model $R_{wp}$ $(\mu\pm\sigma)\downarrow$ WD $(\mu\pm\sigma)\downarrow$
none U-deCIFer $1.24$$\pm 0.26$ $0.49$$\pm 0.31$
deCIFer $0.32$$\pm 0.34$ $0.14$$\pm 0.21$
comp. U-deCIFer $0.82$$\pm 0.41$ $0.28$$\pm 0.29$
deCIFer $0.25$$\pm 0.29$ $0.10$$\pm 0.20$
comp. + s.g. U-deCIFer $0.65$$\pm 0.36$ $0.23$$\pm 0.28$
deCIFer $0.24$$\pm 0.29$ $0.10$$\pm 0.20$
none U-deCIFer 99.82 98.87 94.30 99.47 93.49 0.00
deCIFer 99.42 98.85 93.69 99.46 92.66 5.01
comp. U-deCIFer 99.87 99.09 94.40 99.46 93.78 49.30
deCIFer 99.68 99.21 94.37 99.55 93.73 91.50
comp.+s.g. U-deCIFer 99.85 98.88 94.51 99.47 93.72 87.07
deCIFer 99.74 99.26 94.38 99.58 93.90 94.53
Section 4.2 Robustness to Perturbations in PXRD Conditioning
@@
Table 3 (and Figure 15 in Appendix) show that while additive noise results in slight performance degradation, the model remains robust to in-distribution noise and peak broadening. We also observe moderate performance degradation with OOD broadening or significantly higher noise levels. Unsurprisingly, we observe that lower-symmetry crystal systems remain more challenging to predict under perturbed conditions across all crystal systems. This is depicted in Figure 16 in the Appendix.
Table 3: Performance of deCIFer under varying levels of noise and peak broadening in PXRD inputs for NOMA and CHILI-100K test sets, evaluated with compositional descriptors. We evaluate two in-distribution (ID) scenarios and one out-of-distribution (OOD) scenario for the PXRD input. Reported metrics include residual weighted profile ($R_{\mathrm{wp}}$), Wasserstein distance (WD), overall validity, and match rate.
[…]
NOMA ID: (0.00, 0.05) 0.25$\pm$0.29 0.10$\pm$0.20 93.73 91.50
ID: (0.05, 0.10) 0.31$\pm$0.30 0.12$\pm$0.21 93.77 89.28
OOD: (0.10, 0.20) 0.65$\pm$0.34 0.23$\pm$0.29 91.66 77.66
CHILI-100K ID: (0.00, 0.05) 0.70$\pm$0.37 0.22$\pm$0.21 41.83 37.34
ID: (0.05, 0.10) 0.73$\pm$0.36 0.22$\pm$0.21 40.95 35.97
OOD: (0.10, 0.20) 0.87$\pm$0.33 0.24$\pm$0.20 33.62 26.09
Section 4.3 OOD Evaluation on CHILI-100K
@@
To evaluate the generalization of deCIFer to more complex crystal systems, we tested its performance on the CHILI-100K dataset Friis-Jensen et al. (2024) which had no overlap with the NOMA training data for deCIFer. Unlike synthetic datasets, CHILI-100K presents a closer approximation to real-world challenges, including complex structural motifs and a broader distribution of crystal symmetries. Additionally, CHILI-100K contains a significantly higher proportion of lower-symmetry structures compared to synthetic datasets like NOMA (see the sample distribution in Appendix Figure 17).
@@
The results on CHILI-100k dataset are summarized in Table 3 and Figure 7. These show that deCIFer maintains a reasonable level of structural accuracy on this challenging dataset. The relatively low validity score is partly due to challenges with bond length validity, which was notably lower than the other validity metrics. A full breakdown of the validity metrics for this experiment is available in Section A.10 in the Appendix.
@@
Despite the performance drop, deCIFer demonstrated robustness to added noise and peak broadening in PXRD inputs, with stable $R_{\mathrm{wp}}$ and WD values across perturbed conditions as seen in Figure 7. This stability, along with its ability to generate a variety of structural features, suggests that deCIFer could be useful for practical applications involving experimentally derived PXRD data.
Figure 7: Distribution of $R_{\mathrm{wp}}$ and WD for the CHILI-100K test set, presented as violin plots with overlain boxplots, for deCIFer using both PXRD conditionings and composition (comp.). The distributions show predictions under three PXRD conditioning scenarios: (top) clean, (middle) maximum noise and broadening, and (bottom) an out-of-distribution PXRD scenario.
@@
PXRD-steered structure generation: The experiments on NOMA and CHILI-100k in Section 4 clearly show the utility of including conditioning data such as PXRD when performing CSP. Unlike unconditional structure generation, generating crystal structures that are steered by scattering data such as PXRD is more useful in applications where a specific target property might be desirable.
@@
Consistency in CIF generation: To investigate the consistency and variability of deCIFer, we generated 16K CIFs for the same PXRD profile with different crystal descriptors (“none”, “comp.”, and “comp+s.g.”). Figure 8 illustrates our findings for a challenging monoclinic crystal system of Sr2Cd2Se4 from the NOMA test set. When unconstrained by crystal descriptors, the model generates a wide diversity of cell-parameters, compositions, and space groups, yet the $R_{\mathrm{wp}}$ values tend to cluster in a relatively narrow range. In contrast, imposing compositional and, especially, space group constraints yields much tighter cell-parameter distributions and a broader $R_{\mathrm{wp}}$ distribution. The broader distributions highlight how the $R_{\mathrm{wp}}$ metric is highly sensitive to even small structural deviations and shows the importance of complementing less sensitive metrics like the WD with other validation methods. Despite these variations, the overall structural match rate to the reference CIF remains high. If the composition or space group is known with high confidence, incorporating these constraints can speed up convergence to a more accurate structural solution. Conversely, if the objective is to explore the material landscape more broadly, unconstrained generation with PXRD can be advantageous.
@@
Figure 8: Illustration of deCIFer-sampled structures for a single PXRD profile of the monoclinic crystal system of Sr2Cd2Se4 (16K samples). $R_{\mathrm{wp}}$ comparisons were made with transformation $\tau_{0}$, void of noise and with FWHM$=0.05$, and with three crystal descriptor scenarios. a) Reference structure. b) Empiric probability distribution of generating a CIF with a given $R_{\mathrm{wp}}$ to the reference structure. c) Representative examples of generated structures showing best (lowest $R_{\mathrm{wp}}$), median, worst (highest $R_{\mathrm{wp}}$), and additional diverse examples. d) Distribution of the sampled crystal systems. e) Histograms of cell- lengths (a,b,c) and angles ($\alpha,\beta,\gamma$) for the 16K sampled structures, with reference cell parameters shown as dotted lines.
@@
Limitations: In materials science, the concept of data leakage is nuanced due to the challenges in defining novelty and diversity, as emphasized in recent discussions on scaling AI for materials discovery Cheetham and Seshadri (2024). Although we stratify the NOMA dataset by space group bins and use CHILI-100K exclusively for testing, there remains the possibility of implicit overlaps, such as structurally similar crystal entries or shared compositional biases. However, the inherent redundancy in materials data, where many “new" compounds are variations of existing structures, might simply reflect the natural landscape of materials science. Furthermore, the vast size and diversity of the NOMA dataset as well as the thorough preprocessing steps (e.g., de-duplication, filtering, and standardization), coupled with the independent curation of CHILI-100K, significantly mitigate the impact of any potential overlaps.
We rely on two metrics to validate the performance of deCIFer, $R_{\mathrm{wp}}$ and WD. While $R_{\mathrm{wp}}$ is the standard in Rietveld refinement, and is highly sensitive to small deviations in atomic positions or cell parameters, it can overemphasize noise or strong peaks. Conversely, WD captures broad differences more robustly, but can be disproportionally sensitive to outlier peaks. Each reveal different aspects of pattern agreement, so employing them in tandem, along with comparing structural similarities is necessary for robust validation.
Future Work: This work opens several directions for further exploration and improvement.
Another direction could be to integrate reinforcement learning from human feedback (RLHF) to guide the model more directly toward generating accurate and chemically valid structures Ziegler et al. (2019). By defining a reward function tailored to properties such as low $R_{\mathrm{wp}}$ values, structural integrity, and adherence to chemical constraints, and interaction with a human expert, RLHF could further refine the model’s outputs.
@@
In this work, we introduced deCIFer as a data-informed approach to CSP. deCIFer is an autoregressive transformer-based language model that generates CIFs directly from PXRD data. By conditioning on simulated PXRD profiles, deCIFer captures fine-grained structural information beyond what is possible with composition- or symmetry-based information alone. Evaluations on large synthetic and experimentally derived datasets highlight deCIFer’s robust performance in scenarios with varying levels of noise and peak broadening. deCIFer was developed using lab-scale resources with a single GPU. This holds potential for diverse experiments going forward. The conditional prompting paradigm introduced in deCIFer can be used to integrate other experimental data. This can be essential to accelerating material design processes.

### v3

@@ Abstract
deCIFer has two learnable components, $f_{{\mathbf{\Phi}}}$ and $f_{{\mathbf{\Theta}}}$. The encoder $f_{{\mathbf{\Phi}}}$ is a 2-layer MLP that takes a PXRD profile of dimension $1000$ and outputs a $512$-dimensional embedding. This embedding is prepended as the conditioning-token to the sequence of CIF tokens, where each CIF token is also of $D=512$ dimensions. $f_{{\mathbf{\Theta}}}$ is a decoder-only transformer (Vaswani et al., 2017) with $8$ layers, each containing $8$ attention heads. We use a context length of $3076$ and a batch size of $32$. A linear warm-up of the learning rate is applied over the first $100$ epochs, followed by a cosine decay schedule using AdamW (Loshchilov and Hutter, 2017) ($\beta_{1}=0.9,\beta_{2}=0.95$, weight decay $10^{-1}$) for $50$K epochs with gradient accumulation of $40$ steps on a single NVIDIA A100 GPU with mixed-precision acceleration. All aspects of the model architecture were implemented in Pytorch (Paszke et al., 2019). $f_{{\mathbf{\Phi}}}$ has $\approx\!0.78$M trainable parameters and $f_{{\mathbf{\Theta}}}$ contains $\approx\!26.94$M, resulting in a deCIFer model with $27.72$M parameters. The code is open-source and additional details can be found in Section A.9 in the Appendix.
Evaluation:
Figure 2 summarizes the evaluation pipeline which shows how a reference CIF from the test set is used to generate discrete peaks, $\mathcal{P}=\{(q_{k},i_{k})\}_{k=1}^{n}$, which are transformed into a PXRD profile $\mathbf{y}=\tau_{\mathrm{fixed}}(\mathcal{P})$, where $\tau_{\mathrm{fixed}}$ is selected according to the desired experimental setting. The resulting PXRD, along with any optional crystal descriptors, is tokenized and passed to deCIFer to produce a new CIF ($\text{CIF}^{*}$). For evaluation, the generated structures are compared using three common metrics. Here a clean reference transformation $\tau_{0}$ (FWHM$=0.05$, $\sigma_{\mathrm{noise}^{2}}=0$) is used.
Figure 2: Evaluation pipeline: A test set CIF generates a PXRD profile, tokenized for deCIFer to produce a new CIF, compared to the reference using a clean transformation.
1) Residual weighted profile ($R_{\mathrm{wp}}$) is applied to convolved continuous PXRD profiles, measuring the discrepancy between a reference PXRD profile, ${\mathbf{y}}$, and a generated PXRD profile, ${\mathbf{y}}^{*}$. Formally: $R_{\mathrm{wp}}=\sqrt{\frac{\sum_{i}w_{i}(y_{i}-y_{i}^{*})^{2}}{\sum_{i}w_{i}y_{i}^{2}}}$ with all weights $w_{i}=1$ here, following convention.
2) Match rate (MR) uses StructureMatcher (Ong et al., 2013) to assess structural similarity between reference and generated CIFs. Two structures are considered a match if their lattice parameters, atomic coordinates, and space group symmetries are within the defined tolerances. MR is the fraction of matching structures. See Section A.7 in the Appendix for more information.
3) Validity (Val.) ensures the internal consistency of each generated CIF by checking formula consistency, site multiplicity, bond lengths, and space group alignment. A CIF is deemed overall valid only if it passes all four checks. Detailed explanations of these metrics are provided in Section A.6 in the Appendix.
@@
We conducted a series of experiments to evaluate deCIFer’s ability to perform CSP when conditioned on PXRD profiles, space group, and composition. First, we compare against three recent state-of-the-art CSP models. Next, we compare the baseline performance of deCIFer against an unconditioned variant (U-deCIFer) to assess the direct impact of PXRD conditioning on the generated structures. Next, we introduce controlled noise and peak broadening into the input PXRD data, examining deCIFer’s robustness under more challenging scenarios resembling real-world PXRD data. Finally, we apply deCIFer (trained on NOMA) to CHILI-100K (Friis-Jensen et al., 2024) to demonstrate its scalability and performance on more complex crystal systems. In these experiments, we generate one CIF for each reference sample in the test sets. For consistent evaluation, each reference CIF is also processed through the fixed, clean transformation $\tau_{0}$ (FWHM $=0.05$, $\sigma_{\mathrm{noise}}^{2}=0$).
@@
Baseline Comparisons with State-of-the-art:
To evaluate the structure generation quality of deCIFer, we compare against three recent state-of-the-art CSP models: CDVAE (Xie et al., 2022), DiffCSP (Jiao et al., 2023), and CrystaLLM (Antunes et al., 2024). Following the single-sample evaluation setup used in prior work (Antunes et al., 2024), we generate one structure per composition and assess accuracy using two metrics: match rate and root-mean-square error (RMSE) in atomic positions (Å). To ensure a fair and consistent evaluation, we adopt the official train-test splits provided by DiffCSP (Jiao et al., 2023) for all four benchmark datasets: Perov-5 (Castelli et al., 2012a; b), Carbon-24 (Pickard, 2020), MP-20 (Jain et al., 2013a), and MPTS-52 (Baird, 2023).
As shown in Table 1, deCIFer outperforms all baselines on Perov-5 and Carbon-24, highlighting the benefit of PXRD conditioning when the diffraction data are informative. However, on MP-20 and MPTS-52, CrystaLLM-large outperforms deCIFer, likely due to its exclusive emphasis on composition-structure mappings, which align closely with the test set. This highlights a key trade-off in experimental data informed CSP: while conditioning can enhance structure generation in complex cases, it can also introduce ambiguity when the PXRD signal is weak or the model’s training priors are strong.
Table 1: Performance comparison on four public CSP benchmarks: Perov-5 (Castelli et al., 2012a; b), Carbon-24 (Pickard, 2020), MP-20 (Jain et al., 2013a), and MPTS-52 (Baird, 2023). Following the single-sample evaluation protocol used in prior work (Antunes et al., 2024), one structure is generated per test composition. We report the match rate (%) based on structural equivalence under StructureMatcher and the atomic root-mean-square error (RMSE in Å) after alignment.
Perov-5 Carbon-24 MP-20 MPTS-52
[…]
CDVAE 45.31 0.1138 17.09 0.2969 33.90 0.1045 5.34 0.2106
DiffCSP 52.02 0.0760 17.54 0.2759 51.49 0.0631 12.19 0.1786
CrystaLLM-small 47.95 0.0966 21.13 0.1687 55.85 0.0437 17.47 0.1113
CrystaLLM-large 46.10 0.0953 20.25 0.1761 58.70 0.0408 19.21 0.1110
U-deCIFer 50.55 0.1177 17.33 0.1526 44.99 0.0784 11.80 0.1563
deCIFer 85.29 0.0491 37.16 0.1970 43.51 0.0763 11.44 0.1346
Importance of PXRD Conditioning:
@@
Figure 3 shows that deCIFer achieves lower average $R_{\mathrm{wp}}$ than U-deCIFer across all descriptor settings. Figure 3 visualizes the distributions of $R_{\mathrm{wp}}$ for both models, highlighting the large performance improvement due to PXRD conditioning, and a smaller but still notable improvement due to composition conditioning. Figure 4 shows that structures of more common crystal systems lead to better outcomes, while under-represented, lower symmetry crystal systems remain challenging. The figure further illustrates the effectiveness of PXRD conditioning with three examples from the NOMA test set, showcasing the range of generated structures, from near-perfect alignment to structural mismatch. While U-deCIFer benefits from the addition of crystal descriptors, it never reaches the accuracy of deCIFer. Overall, these results demonstrate that conditioning on PXRD data and crystal descriptors substantially enhances the model’s ability to generate CIFs with close alignment to desired PXRD profiles.
Desc. Model $R_{wp}$ $(\mu\pm\sigma)\downarrow$ Val. (%) $\uparrow$ MR (%) $\uparrow$
none U-deCIFer $1.24$$\pm 0.26$ 93.49 0.00
deCIFer $0.32$$\pm 0.34$ 92.66 5.01
comp. U-deCIFer $0.82$$\pm 0.41$ 93.78 49.30
deCIFer $0.25$$\pm 0.29$ 93.73 91.50
comp.+s.g. U-deCIFer $0.65$$\pm 0.36$ 93.72 87.07
deCIFer $0.24$$\pm 0.29$ 93.90 94.53
Figure 3: Left: Distribution of $R_{\mathrm{wp}}$ for deCIFer and U-deCIFer on the NOMA test set with boxplots. Lower $R_{\mathrm{wp}}$ indicates better CIF alignment. Right: Performance for 20K NOMA test samples using deCIFer and U-deCIFer with different descriptors: none (no descriptors), comp. (composition), and comp.+ s.g. (composition + space group). Metrics include validity (Val.) and match rate (MR).
Figure 4: Left: Average $R_{\mathrm{wp}}$ by crystal system for deCIFer on the NOMA test set shows better performance for common high symmetry systems and higher $R_{\mathrm{wp}}$ for rare low symmetry systems. Right: Examples from the NOMA test set highlight this trend with predicted structures from PXRD and composition maintaining reasonable matches even for low symmetry systems with higher $R_{\mathrm{wp}}$.
Robustness to Perturbations in PXRD Conditioning
@@
Figure 5 (and Figure 16 in Appendix) show that while additive noise results in slight performance degradation, the model remains robust to in-distribution noise and peak broadening. We also observe moderate performance degradation with OOD broadening or significantly higher noise levels. Unsurprisingly, we observe that lower-symmetry crystal systems remain more challenging to predict under perturbed conditions across all crystal systems. This is depicted in Figure 14 in the Appendix.
OOD Evaluation on CHILI-100K
@@
To evaluate the generalization of deCIFer to more complex crystal systems, we tested its performance on the CHILI-100K dataset (Friis-Jensen et al., 2024) which had no overlap with the NOMA training data for deCIFer. Unlike synthetic datasets, CHILI-100K presents a closer approximation to real-world challenges, including complex structural motifs and a broader distribution of crystal symmetries. Additionally, CHILI-100K contains a significantly higher proportion of lower-symmetry structures compared to synthetic datasets like NOMA (see the sample distribution in Appendix Figure 15).
@@
The results on CHILI-100k dataset are summarized in Figure 5. deCIFer maintains a reasonable level of structural accuracy on this challenging dataset. The relatively low validity score is partly due to challenges with bond length validity, which was notably lower than the other validity metrics. A full breakdown of the validity metrics for this experiment is available in Section A.6 in the Appendix.
@@
Despite the performance drop, deCIFer demonstrated robustness to added noise and peak broadening in PXRD inputs, with stable $R_{\mathrm{wp}}$ values across perturbed conditions as seen in Figure 5. This stability, along with its ability to generate a variety of structural features, suggests that deCIFer could be useful for practical applications involving experimentally derived PXRD data.
Table 2: Performance on CHILI-100K without (U-deCIFer) and with (deCIFer) PXRD conditioning.
Model $R_{\mathrm{wp}}\downarrow$ Val. (%$\uparrow$ MR (%$\uparrow$
U-deCIFer $0.96\pm 0.32$ $43.26$ $25.92$
deCIFer $0.70\pm 0.37$ $41.83$ $37.34$
Dataset ($\sigma_{\mathrm{noise}}^{2}$, FWHM) $R_{\mathrm{wp}}\;(\mu\pm\sigma)\downarrow$ Val. (%)\;$\uparrow$ MR (%)\;$\uparrow$
NOMA ID: (0.00, 0.05) 0.25$\pm$0.29 93.73 91.50
ID: (0.05, 0.10) 0.31$\pm$0.30 93.77 89.28
OOD: (0.10, 0.20) 0.65$\pm$0.34 91.66 77.66
CHILI-100K ID: (0.00, 0.05) 0.70$\pm$0.37 41.83 37.34
ID: (0.05, 0.10) 0.73$\pm$0.36 40.95 35.97
OOD: (0.10, 0.20) 0.87$\pm$0.33 33.62 26.09
Figure 5: Left: Distribution of $R_{\mathrm{wp}}$ for deCIFer on NOMA and CHILI-100K using PXRD conditioning and composition (comp.) across three scenarios: clean, high noise/broadening, and out-of-distribution (OOD), with noise and FWHM values indicated. Right: Corresponding table of performance metrics, including $R_{\mathrm{wp}}$, overall validity (Val.), and match rate (MR) for each scenario.
To assess the impact of PXRD conditioning on OOD data, we compare deCIFer with U-deCIFer on the CHILI-100K test set. Results are shown in Table 2. We observe that PXRD conditioning lowers $R_{\mathrm{wp}}$ and has a significant positive effect on the match rate, showcasing that the conditioning drives improvement and not merely compositional priors.
@@
PXRD-driven Structure Generation: The experiments on NOMA and CHILI-100K in Section 5 clearly show the utility of including conditioning data such as PXRD when performing CSP. deCIFer consistently achieves strong performance on the NOMA test set, where the diversity of crystal systems and the availability of high-quality PXRD signals allow the model to accurately generate structures that align with the diffraction data. This highlights a key advantage of PXRD conditioning: when the diffraction data are clear and informative, deCIFer can use them to resolve complex atomic arrangements, even in cases where composition-only models struggle.
However, the results also make it clear that deCIFer, like any PXRD-conditioned model, cannot be used as a black-box end-to-end solver. The model’s ability to generate accurate structures is tied to the quality and relevance of the PXRD signal. On datasets like MP-20, where learned composition–structure priors dominate, PXRD conditioning can sometimes interfere with accurate generation by introducing conflicting guidance. This tension between learned priors and PXRD conditioning is not unique to deCIFer but reflects a broader challenge in PXRD-informed CSP.
@@
Consistency in CIF generation: To further investigate the consistency and variability of deCIFer, we generated 16K CIFs for the same PXRD profile with different crystal descriptors (“none”, “comp.”, and “comp+s.g.”). Figure 6 illustrates our findings for a challenging monoclinic crystal system of Sr2Cd2Se4 from the NOMA test set. When unconstrained by crystal descriptors, the model generates a wide diversity of cell-parameters, compositions, and space groups, yet the $R_{\mathrm{wp}}$ values tend to cluster in a relatively narrow range. In contrast, imposing compositional and, especially, space group constraints yields much tighter cell-parameter distributions and a broader $R_{\mathrm{wp}}$ distribution. The broader distributions highlight how the $R_{\mathrm{wp}}$ metric is highly sensitive to even small structural deviations and shows the importance of complementing with other validation methods. Despite these variations, the overall structural match rate to the reference CIF remains high. If the composition or space group is known with high confidence, incorporating these constraints can speed up convergence to a more accurate structural solution. Conversely, if the objective is to explore the material landscape more broadly, unconstrained generation with PXRD can be advantageous.
@@
Figure 6: deCIFer-sampled structures for a monoclinic Sr2Cd2Se4444 PXRD profile (16K samples). a) Reference structure. b) Distribution of $R{\mathrm{wp}}$ for generated CIFs. c) Examples of generated structures showing best, median, and diverse samples. d) Distribution of sampled crystal systems. e) Histograms of cell lengths (a, b, c) and angles ($\alpha,\beta,\gamma$) with reference values as dotted lines.
@@
Limitations: In materials science, the concept of data leakage is nuanced due to the challenges in defining novelty and diversity, as emphasized in recent discussions on scaling AI for materials discovery (Cheetham and Seshadri, 2024). Although we stratify the NOMA dataset by space group bins and use CHILI-100K exclusively for testing, there remains the possibility of implicit overlaps, such as structurally similar crystal entries or shared compositional biases.
Furthermore, the vast size and diversity of the NOMA dataset as well as the thorough preprocessing steps (e.g., de-duplication, filtering, and standardization), coupled with the independent curation of CHILI-100K, significantly mitigate the impact of any potential overlaps.
While PXRD conditioning enhances structure prediction by introducing rich physical constraints, it also inherits a well-known limitation: the inability to distinguish between homometric structures. These are distinct atomic arrangements that yield indistinguishable diffraction patterns (Patterson, 1944). This ambiguity has been documented in real structural analyses (Schneider et al., 2010). Because metrics such as $R_{\mathrm{wp}}$ operate only on the diffraction pattern, they cannot capture homometric degeneracy. Nevertheless, several strategies can help mitigate this limitation (Shen et al., 2022). Particularly, our experiments with deCIFer show that even partial inclusion of complementary information, such as composition, can mitigate near-degeneracy cases where the only differences in the PXRD are subtle variations in relative peak intensities.
@@
In this work, we introduced deCIFer as a data-informed approach to CSP. deCIFer is an autoregressive transformer-based language model that generates CIFs directly from PXRD data developed using lab-scale compute resources with a single GPU. By conditioning on simulated PXRD profiles, deCIFer captures fine-grained structural information beyond what is possible with composition- or symmetry-based information alone. Evaluations on large synthetic and experimentally derived datasets highlight deCIFer’s robust performance in scenarios with varying levels of noise and peak broadening.
Ultimately, deCIFer represents a step forward in PXRD-informed CSP, bridging the gap between purely composition-based generative models and structure generation directly guided by experimental data. It is most useful when the diffraction signal is informative and when combined with careful downstream verification. For real-world use, this balance must be clearly understood, and model predictions must be treated as high-quality starting points rather than definitive solutions.